OpenStudy (anonymous):

Please help! If F ( x) = 4 - 2 x, find the ordered pairs corresponding to the following number in the domain: 0 The ordered pair is ?

OpenStudy (anonymous):

If the domain =0 i.e. x= 0 then f(x) = 4-2x => F(0) = 4-0=4 therefore the ordered pair could be: (0,4)

OpenStudy (anonymous):

Domain = x values Range = F(x) values an ordered pair would be [ x, F(x)] = [ 0, 4]

OpenStudy (anonymous):

Thank you so much for your help chemwile!

OpenStudy (anonymous):

vw